# SKATE, a Scalable Tournament Eval: Weaker LLMs differentiate between stronger ones using verifiable challenges

### v1

@@ Abstract
Evaluating the capabilities and risks of foundation models is paramount, yet current methods demand extensive domain expertise, hindering their scalability as these models rapidly evolve. We introduce SKATE: a novel evaluation framework in which large language models (LLMs) compete by generating and solving verifiable tasks for one another. Our core insight is to treat evaluation as a game: models act as both task-setters and solvers, incentivized to create questions which highlight their own strengths while exposing others’ weaknesses. SKATE offers several key advantages, balancing scalability, open-endedness, and objectivity. It is fully automated, data-free, and scalable, requiring no human input or domain expertise. By using verifiable tasks rather than LLM judges, scoring is objective. Unlike domain-limited programmatically-generated benchmarks (e.g. chess-playing or spatial reasoning), having LLMs creatively pose challenges enables open-ended and scalable evaluation. As a proof of concept, we introduce LLM-set code-output-prediction (COP) challenges as a verifiable and extensible framework in which to test our approach. Using a TrueSkill-based ranking system, we evaluate six frontier LLMs and find that: (1) weaker models can reliably differentiate and score stronger ones, (2) LLM-based systems are capable of self-preferencing behavior, generating questions that align with their own capabilities, and (3) SKATE automatically surfaces fine-grained capability differences between models. Our findings are an important step towards general, scalable evaluation frameworks which can keep pace with LLM progress.
@@ 1 Introduction
The extraordinary evolution of large language models (LLMs) underscores the need for robust, scalable, and unbiased evaluation methodologies. As these models are increasingly deployed in high-stakes settings - such as healthcare, education, legal reasoning, and autonomous decision-making - the risks of mistake or misuse become more pronounced (Weidinger et al. 2021; Bommasani et al. 2022; Perez et al. 2022a). Current evaluation frameworks, however, often fall short of these goals. They typically require extensive human expertise or qualitative assessments. Furthermore, they are often costly to develop and maintain, not scalable enough to keep pace with the advancements of new models, and can be manipulable, static, or gameable (Weng 2024; METR 2025).
@@
LLMs setting tasks for themselves inherently presents limitations, such as potential self-preferencing biases (Gera et al. 2025) or awareness of being tested (Perez et al. 2022b). To mitigate these issues, our framework is designed as a peer-challenge game where LLMs actively compete, creating a more objective and dynamic evaluation environment.
@@
As a proof of concept, we introduce LLM-set Code-Output-Prediction (COP) challenges. This verifiable and extensible framework serves as a concrete testbed for our approach. We apply our peer-challenge game to six frontier LLMs, employing a TrueSkill (Herbrich, Minka, and Graepel 2006)-based ranking system to quantitatively assess their performance. While this paper focuses on COP, superficially a narrow task, we claim that this approach has broad applicability: since computational tasks can be expressed as code execution problems, COP provides a general substrate for evaluating model capabilities across diverse domains.
@@
    Evidence that weaker models can reliably score stronger ones, with transitive and stable comparisons.
@@
    Evidence that LLM-based systems exhibit a capacity to design self-preferencing questions.
@@
    Automatic discovery of differentiating questions which expose capability differences between models.
@@ 2 Related work
Many benchmarks aim to evaluate the capabilities of AI models. Some are general purpose (Hendrycks et al. 2021a) or for general agentic tasks (Mialon et al. 2023), others for particular skills or knowledge areas (Li et al. 2024; Hendrycks et al. 2021b). Evaluations of coding performance range from single-completion generation of code (Chen et al. 2021), to agentic resolution of GitHub issues (Jimenez et al. 2024), to ML engineering tasks (Huang et al. 2024), and to many others. Criticism includes suggestions that benchmarks may be unrealistic (Kapoor et al. 2024; Becker et al. 2025), or distract from higher-priority safety interventions (Ren et al. 2024).
Evaluations often take significant effort to produce (Phan et al. 2025)), and those which are tuned to be sensitive to model performance at the point of publication are often quickly "saturated", as models progress faster than expected (Bengio et al. 2024) and reach indistinguishably-high performance (Hendrycks et al. 2021b). Our paper describes an approach where the generation of the evaluation is automated, and where a cohort of LLM peers can distinguish model capabilities, scaling as ever more capable models are developed.
@@
Arbitrary tasks can be scored using LLM judges, but expose the evaluation to the judges’ biases (Gera et al. 2025). Game-based evaluations provide advantages including objective outcomes, verifiable moves, reduced data contamination (representation in training data), and scalability. Even simple grid-based games such as Tic-Tac-Toe and Connect Four (Topsakal, Edell, and Harper 2024) reveal variations in LLM performance across different games and prompt types. Chess has also been used as a testbed (Zhang et al. 2025; Wang et al. 2024; Diallo and Akhloufi 2025), including a leaderboard against random play (Saplin 2024), showing that both fine-tuning and language explanations can enhance the performance and reasoning capability of large language models, as can tools such as RAG. In all settings, the ultimate goal is to find strategic planning and decision-making abilities which generalize to real-world tasks.
@@
The idea of LLMs generating tasks has been explored in prior work. Lu, Hu, and Clune (2025) focus on high-level capabilities and qualitative comparisons between models; crucially, their tasks are open-ended and validated using LLMs-as-judges. We propose to generate objective, automatically verifiable tasks, and provide a scalable mechanism to use these tasks to rank models.
Perez et al. (2022b) have used LLM-generated datasets to automatically probe a broad range of nuanced properties such as sycophancy, and concerning goals such as resource acquisition and goal preservation. However, while the authors find that crowdworkers agree with $90\mbox{--}100\%$ of labels, the questions set by LLMs are not automatically verifiable and it is unclear whether the dataset-generation approach fully avoids the pitfalls of LLM-judges (Gera et al. 2025).
Other work generates tasks with templates (Weston et al. 2015) or programmatically (Johnson et al. 2016), which, while verifiable, limit the diversity, expressivity and scalability seen in our model-written COP approach.
@@
Automated Red Teaming (ART) finds cases where a target LLM behaves in a harmful way, by generating test cases (“red teaming”) using another LLM (Perez et al. 2022a). Active research has led to many advances, including curiosity-driven exploration (Hong et al. 2024), structured search (Mehrotra et al. 2024), reinforcement learning, prompt engineering and optimization, and transferability (Raheja, Pochhi, and Curie 2024). However, to our knowledge these approaches focus exclusively on jail-breaking, rather than an open-ended exploration of the limits of a model’s capabilities.
@@
Concerned by the potential future need to supervise systems which broadly outperform humans, and which are hard to evaluate, various groups have researched methods for weaker models to elicit the capabilities of stronger ones (Burns et al. 2023; Bowman et al. 2022). While these methods constitute progress towards aligning superhuman models to the goals of weaker models, with SKATE weaker models can evaluate the capabilities of stronger ones, to choose which strong student to teach.
3 Background
@@
The idea of using LLMs to evaluate other LLMs is appealing, but unconstrained question setting often yields questions that are subjective or costly to verify (e.g. “write a moving poem", or complex legal queries). Whilst expensive verification can be mitigated by “LLMs-as-judges", this approach risks inherent biases (Gera et al. 2025). To counter this, we design our framework around verifiable tasks - those with clear, systematic, and objective assessable solutions. For the purposes of our work, we choose Code-Output-Prediction (COP) tasks as our testbed: given a block of code, what does the code output? Correct answers can be determined by via a code-execution sandbox. Other examples of verifiable tasks include games, writing code to pass unit tests, and factual questions with definitive answers. We bound the output space of model answers, to quantify a model’s performance on a question given the probabilistic nature of LLMs, by making our tasks multiple-choice. "Distractor" incorrect options are also generated by the question-asking model, while the correct answer is determined using a code-execution sandbox. While our experiments focus on COP, the SKATE framework is general-purpose and can be instantiated with any such verifiable task type.
@@
LLM responses to multiple choice questions (MCQs) are sensitive to factors like option ordering (Zheng et al. 2024) and option content, see Figure. 8. To account for this, we estimate a model’s score on a question by sampling it multiple times with randomly permuted answer sets. Each includes the correct answer and three distractors selected from a pool of nine. We repeat this process until the estimated accuracy, denoted p(correct), has standard deviation below a stability threshold ($\sigma^{*}=0.05$). This yields robust per-question probabilities $\in[0,1]$ that we use to compute both relative and absolute rankings of models (see Section 4.3 and Appendix A).
@@
Measuring the similarity of generated tasks is crucial for categorizing capabilities and preventing models from “reward hacking" through repeated, similar questions. We aim to incentivise exploration of diverse capabilities. To this end, we use vector embeddings (specifically openai-text-embedding-3-small (Neelakantan et al. 2022)) and define question similarity based on a cosine similarity distance threshold. From empirical analysis of a diverse dataset of 2,977 questions, we found a threshold on $d_{ij}=1-d_{ij}^{\cos}$ of $d_{\text{thresh}}=0.336$ effectively groups distinct questions. More information is in Appendix B.
@@ 4 Peer-Challenge LLM Benchmark Game
In this section we define the format of our “Game of SKATE", and how it is scored. With this in place, we define various “augmentation strategies” which allow models to take advantage of various pieces of information (e.g. models’ p(correct) on previous questions).
Most evaluation frameworks treat models as passive subjects: evaluated via curated benchmarks or third-party judges. In contrast, SKATE frames evaluation as an adversarial game: models actively participate as both task designers and solvers.
@@ 4.1 Defining The Game
It is possible to give all players access to full information about the game state in each round, including all previous questions asked by all players (including themselves), and the p(correct) scores achieved by all players on those questions. However, since this represents a large amount of information, in Section C we propose a variety of augmentation strategies which permit the task-setters to ingest increasing amounts of this information in the context window during task setting. Concretely: we draw a distinction between the game rules which in principle allow full access to game state information, and augmentation strategies which involve players using some, or all, of this information to help them play optimally.
@@ 4.2 Incentives
At the level of the prompt, we tell the task-setting LLMs that they are playing a game and will be rewarded as follows (see Appendix F for the full prompt):
  • +1 if it successfully created a valid question.
  • +1 for each question it gets correct (for both its own q and those of its competitors).
We emphasize that these rewards are not the scoring mechanism we use to rank the models (which we detail in Section 4.3). but rather information used only in the prompt, designed to incentivize the following behavior of the task-setters:
  • Writing valid questions.
  • Writing questions which they can answer.
  • Writing questions which their opponents get wrong.
One way of re-phrasing this is in terms of discriminatory questions, defined as valid questions which the task-setter answers correctly whilst the competitors answer incorrectly. Our game is designed such that optimal behavior for a player is to write 50 discriminatory questions. By requiring suitably unique questions, we require the players to isolate as many discriminatory niches as possible; that is, to maximise the number of distinct areas in which they excel and their competitors do not.
@@ 4.3 Ranking
We use a ranking system to order models after the game. We employ TrueSkill (Herbrich, Minka, and Graepel 2006), which is a Bayesian ranking system developed by Microsoft for Xbox Live. It is designed to rate the skill of players in a competitive game whilst taking into account uncertainty of their skill. Instead of a single skill score, TrueSkill represents a player’s skill as a probability distribution, characterized by a mean $\mu$ and a variance $\sigma^{2}$.
@@ 5 Augmentation Strategies
Figure 2: TrueSkill $(\mu,\sigma)$ for four models after one iteration of our game. We compare five augmentation strategies, and score each game in two ways (in the top row using relative TrueSkill, and the bottom row using absolute TrueSkill). We note that relative TrueSkill consistently produces clearer separations between models than absolute scoring, reflecting its higher information content. The values quoted here are the average values over the final 100 update steps.
We propose that sufficiently powerful LLMs could use game data (e.g. questions, and p(correct) values) to optimally construct future questions which favor themselves and disadvantage other players. To test this, we design five “augmentation strategies" which can be viewed as filters which permit types of game-state information to enter the task setters’ contexts. These strategies range from "no-info" (no context about previous questions or performance), up to "full-context" (full information about all previous questions and all players’ p(correct) values on those questions). These five strategies are described in detail in Appendix C and Table 1.
@@
However, the augmentation strategies do influence how models generate questions over time. We show in Appendix C that some strategies lead to more adaptive question setting, with models refining difficulty to match their own abilities and better challenge others. This suggests that the models tested can make limited use of game context information, but do not yet exhibit strong strategic exploitation of that information to be maximally adversarial.
To strike a balance between informativeness and computational cost, we adopt the historical performance augmentation strategy for the rest of our experiments. We expect that as LLM capabilities improve, models will better leverage full-context augmentation to construct highly targeted, adversarial tasks. All experiments use temperature=0.7=0.7.

### v2

@@ Abstract
Evaluating the capabilities and risks of frontier AI models is paramount, yet current methods demand extensive domain expertise, hindering their scalability as these models rapidly evolve. We introduce SKATE: a novel evaluation framework in which large language models (LLMs) compete by generating and solving verifiable tasks for one another. Our core insight is to treat evaluation as a game: models act as both task-setters and solvers, incentivized to create questions which highlight their own strengths while exposing others’ weaknesses. SKATE offers several key advantages, balancing scalability, open-endedness, and objectivity. It is fully automated, data-free, and scalable, requiring no human input or domain expertise. By using verifiable tasks rather than LLM judges, scoring is objective. Unlike domain-limited programmatically-generated benchmarks (e.g. chess-playing or spatial reasoning), having LLMs creatively pose challenges enables open-ended and scalable evaluation. As a proof of concept, we introduce LLM-set code-output-prediction (COP) challenges as a verifiable and extensible framework in which to test our approach. Using a TrueSkill-based ranking system, we evaluate six frontier LLMs and find that: (1) weaker models can score stronger ones consistently, reliably differentiating between them, and (2) LLM-based systems are capable of self-preferencing behavior, generating questions that align with their own capabilities, and (3) SKATE automatically surfaces fine-grained capability differences between models. Our findings are an important step towards general, scalable evaluation frameworks which can keep pace with LLM progress.
@@ 1 Introduction
The extraordinary evolution of large language models (LLMs) underscores the need for robust, scalable, and unbiased evaluation methodologies. As these models are increasingly deployed in high-stakes settings - such as healthcare, education, legal reasoning, and autonomous decision-making - the risks of mistake or misuse become more pronounced (Weidinger et al., 2021; Bommasani et al., 2022; Perez et al., 2022a). Current evaluation frameworks, however, often fall short of these goals. They typically require extensive human expertise or qualitative assessments. Furthermore, they are often costly to develop and maintain, not scalable enough to keep pace with the advancements of new models, and can be manipulable, static, or gameable (Weng, 2024; METR, 2025).
@@
LLMs setting tasks for themselves inherently presents limitations, such as potential self-preferencing biases (Gera et al., 2025) or awareness of being tested (Perez et al., 2022b). To mitigate these issues, our framework is designed as a peer-challenge game where LLMs actively compete, creating a more objective and dynamic evaluation environment.
@@
As a proof of concept, we introduce LLM-set Code-Output-Prediction (COP) challenges. This verifiable and extensible framework serves as a concrete testbed for our approach. While this paper focuses on COP, superficially a narrow task, we claim that this approach has broad applicability: since computational tasks can be expressed as code execution problems, COP provides a general substrate for evaluating model capabilities, such as reasoning, across diverse domains.
We apply our peer-challenge game to six frontier LLMs, employing a TrueSkill (Herbrich et al., 2006)-based ranking system to quantitatively assess their performance.
@@
    Evidence that weaker models can reliably score stronger ones.
@@
    Evidence that LLM-based systems exhibit a capacity to write self-favoring questions.
@@
    Automatic discovery of differentiating questions, exposing fine-grained capability differences between models.
@@ 2 Related work
Many benchmarks aim to evaluate the capabilities of AI models. Some are general purpose (Hendrycks et al., 2021a) or for general agentic tasks (Mialon et al., 2023), others for particular skills or knowledge areas (Li et al., 2024; Hendrycks et al., 2021b). Evaluations of coding performance range from single-completion generation of code (Chen et al., 2021), to agentic resolution of GitHub issues (Jimenez et al., 2024), to ML engineering tasks (Huang et al., 2024), and to many others. Criticism includes suggestions that benchmarks may be unrealistic (Kapoor et al., 2024; Becker et al., 2025), or distract from higher-priority safety interventions (Ren et al., 2024).
Evaluations often take significant effort to produce (Phan et al., 2025)), and those which are tuned to be sensitive to model performance at the point of publication are often quickly "saturated", as models progress faster than expected (Bengio et al., 2024) and reach indistinguishably-high performance (Hendrycks et al., 2021b). Our paper describes an approach where the generation of the evaluation is automated, and where a cohort of LLM peers can distinguish model capabilities, scaling as ever more capable models are developed.
@@
Arbitrary tasks can be scored using LLM judges, but expose the evaluation to the judges’ biases (Gera et al., 2025). Game-based evaluations provide advantages including objective outcomes, verifiable moves, reduced data contamination (representation in training data), and scalability. Even simple grid-based games such as Tic-Tac-Toe and Connect Four (Topsakal et al., 2024) reveal variations in LLM performance across different games and prompt types. Chess has also been used as a testbed (Zhang et al., 2025; Wang et al., 2024a; Diallo and Akhloufi, 2025), including a leaderboard against random play (Saplin, 2024), showing that both fine-tuning and language explanations can enhance the performance and reasoning capability of large language models, as can tools such as RAG. In all settings, the ultimate goal is to find strategic planning and decision-making abilities which generalize to real-world tasks.
There is existing work involving evaluating LLMs in head-to-head style games. In Duan et al. (2024) LLMs are compared via their head-to-head performance on an environment containing 10 games. In Alyahya et al. (2025) LLMs are evaluated against one another in a competition-based framework with 7 static types of game. An important distinction between this literature and our work is that both papers’ tasks are all human-crafted, whereas in SKATE, LLMs themselves design and pose all the questions in a scalable and dynamic setting.
@@
The idea of LLMs generating tasks has been explored in prior work. Lu et al. (2025) focus on high-level capabilities and qualitative comparisons between models; crucially, their tasks are open-ended and validated using LLMs-as-judges. We propose to generate objective, automatically verifiable tasks, and provide a scalable mechanism to use these tasks to rank models. Perez et al. (2022b) have used LLM-generated datasets to automatically probe a broad range of nuanced properties such as sycophancy, and concerning goals such as resource acquisition and goal preservation. However, while the authors find that crowdworkers agree with $90\mbox{--}100\%$ of labels, the questions set by LLMs are not automatically verifiable and it is unclear whether the dataset-generation approach fully avoids the pitfalls of LLM-judges Gera et al. (2025). Other work generates tasks with templates (Weston et al., 2015) or programmatically (Johnson et al., 2016), which, while verifiable, limit the diversity, expressivity and scalability seen in our model-written COP approach.
@@
Automated Red Teaming (ART) finds cases where a target LLM behaves in a harmful way, by generating test cases (“red teaming”) using another LLM (Perez et al., 2022a). Active research has led to many advances, including curiosity-driven exploration (Hong et al., 2024), structured search (Mehrotra et al., 2024), reinforcement learning, prompt engineering and optimization, and transferability (Raheja et al., 2024). However, to our knowledge these approaches focus exclusively on jail-breaking, rather than an open-ended exploration of the limits of a model’s capabilities.
@@
Concerned by the potential future need to supervise systems which broadly outperform humans, and which are hard to evaluate, various groups have researched methods for weaker models to elicit the capabilities of stronger ones (Burns et al., 2023; Bowman et al., 2022). While these methods constitute progress towards aligning superhuman models to the goals of weaker models, with SKATE weaker models can evaluate the capabilities of stronger ones, to choose which strong student to teach.
3 Setting and scoring Verifiable Tasks
@@
The idea of using LLMs to evaluate other LLMs is appealing, but unconstrained question-setting often yields questions that are subjective or costly to verify (e.g. “write a moving poem", or complex legal queries). Whilst expensive verification can be mitigated by “LLMs-as-judges", this approach risks inherent biases (Gera et al., 2025). To counter this, we design our framework around verifiable tasks - those with clear, systematic, and objective assessable solutions.
Code-Output Prediction as a General Testbed
For the purposes of our work, we experiment using Code-Output-Prediction (COP) tasks: given a block of code, what does the code output? Correct answers can be determined by via a code-execution sandbox. In Appendix C we show that many types of cognitive tasks, such as “counting the number of ‘r’s in “strawberry”, mathematical and spatial reasoning, and various games e.g. chess, can be converted into COP tasks.
Other examples of verifiable tasks include games, writing code to pass unit tests, and factual questions with definitive answers (see Appendix B). For SKATE to be able to use a verifiable task, it must be practical to automatically determine that task’s ground truth answer. In this work, we choose to restrict tasks to be multiple-choice but emphasize that the framework does not require this [^1] In Appendix B we give some examples of different types of COP which are unsuitable for multiple-choice: where automatically generating ground-truth answers is difficult but automatically verifying solutions is trivial (e.g. theorem proving in a formal language). . “Distractor" incorrect options are also generated by the question-asking model, while the correct answer is determined using a code-execution sandbox. While our experiments focus on COP, the SKATE framework is general-purpose and can be instantiated with any such verifiable task type. We note, however, that restricting to a multiple-choice formulation in our experiments inherently restricts the evaluation space, for example by excluding open-ended responses, which may capture different dimensions of model capability.
@@
LLM responses to multiple choice questions (MCQs) are sensitive to factors like option ordering (Zheng et al., 2024) and option content, see Figure 6. To account for this, we estimate a model’s score on a question by sampling it multiple times with randomly permuted answer sets. Each includes the correct answer and three distractors selected from a pool of nine. We repeat this process until the estimated accuracy, denoted p(correct), has standard deviation below a stability threshold ($\sigma^{*}=0.05$). The choice of $\sigma^{*}$ reflects a balance: too large a value would make model comparisons uninformative, while too small a value would be computationally costly. This yields robust per-question probabilities $\in[0,1]$ that we use to compute both relative and absolute rankings of models (see Section 4.3 and Appendix A).
@@
Measuring the similarity of generated tasks is crucial for categorizing capabilities and preventing models from “reward hacking" through repeated, similar questions. We aim to incentivise exploration of diverse capabilities. To this end, we use vector embeddings (specifically openai-text-embedding-3-small (Neelakantan et al., 2022)) [^2] It is possible that using different embedding models could lead to different filtering decisions, but we expect the overall impact of this to be minimal for current models, especially since we observe empirically that the filtering step mostly removes fairly obvious near-duplicates. and define question similarity based on a cosine similarity distance threshold. From empirical analysis of a diverse dataset of 2,977 questions, we found a threshold on $d_{ij}=1-d_{ij}^{\cos}$ of $d_{\text{thresh}}=0.336$ effectively groups distinct questions  [^3] This threshold is a design choice: from empirical observation it separates questions into distinct clusters amongst the models we tested, but may need adjustment for newer models.. More information is in Appendix D.
@@ 4 Peer-Challenge LLM Benchmark Game
In this section we define the format of our “Game of SKATE", and how it is scored. With this in place, we define various “augmentation strategies” which allow models to take advantage of various pieces of information (e.g. models’ p(correct) on previous questions). Most evaluation frameworks treat models as passive subjects: evaluated via curated benchmarks or third-party judges. In contrast, SKATE frames evaluation as an adversarial game: models actively participate as both task designers and solvers.
@@ 4.1 Defining The Game
The uniqueness criterion forces the Asking model to prioritize question variability and avoids the risk of them fixating on one particular capability or question style.
@@
It is possible to give all players access to full information about the game state in each round, including all previous questions asked by all players (including themselves), and the p(correct) scores achieved by all players on those questions. However, since this represents a large amount of information, in Section E we propose a variety of augmentation strategies which permit the task-setters to ingest increasing amounts of this information in the context window during task setting. Concretely: we draw a distinction between the game rules which in principle allow full access to game state information, and augmentation strategies which involve players using some, or all, of this information to help them play optimally.
@@ 4.2 Incentives
At the level of the prompt, we tell the task-setting LLMs that they are playing a game and will be rewarded as follows (see Appendix J for the full prompt): a) +1 if it successfully created a valid question and, b) +1 for each question it gets correct (for both its own q and those of its competitors). We emphasize that these rewards are not the scoring mechanism we use to rank the models (which we detail in Section 4.3). but rather information used only in the prompt, designed to incentivize task-setters to write valid questions which they can answer but their opponents cannot.
We refer to these optimal questions as discriminatory questions, defined as valid questions which the task-setter answers correctly whilst the competitors answer incorrectly. Our game is designed such that optimal behavior for a player is to write 50 discriminatory questions. By requiring suitably unique questions using the similarity metric, we require the players to isolate as many discriminatory niches as possible; that is, to maximise the number of distinct areas in which they excel and their competitors do not.
@@ 4.3 Ranking
We use a ranking system to order models after the game. We employ TrueSkill (Herbrich et al., 2006), which is a Bayesian ranking system designed to rate the skill of players in a competitive game whilst taking into account uncertainty of their skill. Instead of a single skill score, TrueSkill represents a player’s skill as a probability distribution, characterized by a mean $\mu$ and a variance $\sigma^{2}$.
@@ 5 Augmentation Strategies
Figure 2: TrueSkill $(\mu,\sigma)$ for four models after one iteration of our game. We compare five augmentation strategies, and score each game using relative TrueSkill (top row) and absolute TrueSkill (bottom row). Relative TrueSkill produces clearer separation between models, reflecting its higher information content. The values quoted are the average values over the final 100 update steps.
We propose that sufficiently powerful LLMs could use game data (e.g. questions, and p(correct) values) to optimally construct future questions which favor themselves and disadvantage other players. To test this, we design five “augmentation strategies" which can be viewed as filters which permit types of game-state information to enter the task setters’ contexts. These strategies range from "no-info" (no context about previous questions or performance), up to "full-context" (full information about all previous questions and all players’ p(correct) values on those questions). These five strategies are described in detail in Appendix E and Table 1.
@@
However, the augmentation strategies do influence how models generate questions over time. We show in Appendix E that some strategies lead to more adaptive question setting, with models refining difficulty to match their own abilities and better challenge others. This suggests that the models tested can make limited use of game context information, but do not yet exhibit strong strategic exploitation of that information to be maximally adversarial.
To strike a balance between informativeness and computational cost, we adopt the historical performance augmentation strategy for the rest of our experiments. Whilst we expect that as LLM capabilities improve, models will better leverage full-context augmentation to construct highly targeted, adversarial tasks, our experiments cannot formally guarantee this. All experiments use temperature=0.7=0.7.